OpenStudy (azionne):

4(2x -3) + 8(x - 4) = 2(2x + 6) will medal

OpenStudy (azionne):

r u replying ?

OpenStudy (cal.lavender):

8x-12+8x-32=4x+12 16-44=4x+12 28=4x+12 28-12=4x 16=4x 16 divided by 4=x 4=x

OpenStudy (azionne):

gtg be back in 45

OpenStudy (anonymous):

16x-44=4x+12 12x=56 x=?